# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v3

@@ VII-C Embodied Control
Although embodied AI encompasses high-level algorithms, models, and planning modules, its most fundamental and essential component is embodied control. Therefore, it is imperative to consider how to control physical entities and endow them with physical intelligence. Embodied control is closely related to hardware, such as controlling joint movements, end-effector positions, and walking speeds. For robotic arms, knowing the end-effector’s position, how to plan joint trajectories to move the arm to the target? For humanoid robots, knowing the motion patterns, how to control the joints to achieve the target posture? These are critical issues that need to be addressed in control. Several works focus on robotic control, enhancing the flexibility of robotic actions. [387] proposed a vision-based full-body control framework. By connecting a robotic arm and a robotic dog, utilizing all degrees of freedom (12 joints in the legs, 6 joints in the arm, and 1 in the gripper), it tracks the robot dog’s speed and the robotic arm’s end-effector position, achieving more flexible control. Some works [388, 389] employ traditional methods to control bipedal robot walking. MIT’s Cheetah 3 [390], ANYmal [391], and Atlas [392] use robust walking controllers to manage the robots. These developed robots can be used for more agile motion tasks, such as jumping or overcoming various obstacles [393, 394, 395, 396, 397]. Other works [398, 399] focus on the control of humanoid robots, enabling them to perform various actions like humans and mimic human behaviors. Fig .18 illustrates some examples the robotic control.
@@ VIII Challenges and Future Directions
For embodied tasks, it is necessary to introduce causal representation learning. This involves establishing spatial-temporal causal relations across modalities through interactive instructions and state predictions, forming a representation learning system based on interaction and deduction. Moreover, agents need to understand the affordances of objects to achieve adaptive task planning and long-distance autonomous navigation in dynamic scenes. To optimize decision-making, it is necessary to combine counterfactual and causal intervention strategies [403, 404, 405, 406], trace causality from counterfactual and causal intervention perspectives, reduce exploration iterations, and optimize decisions. Constructing a causal graph based on world knowledge and driving sim-to-real transfer of agents through active causal reasoning will form a unified framework for embodied perception, reasoning, and interaction.

### v4

@@ VII-C Embodied Control
Although embodied AI encompasses high-level algorithms, models, and planning modules, its most fundamental and essential component is embodied control. Therefore, it is imperative to consider how to control physical entities and endow them with physical intelligence. Embodied control is closely related to hardware, such as controlling joint movements, end-effector positions, and walking speeds. For robotic arms, knowing the end-effector’s position, how to plan joint trajectories to move the arm to the target? For humanoid robots, knowing the motion patterns, how to control the joints to achieve the target posture? These are critical issues that need to be addressed in control. Several works focus on robotic control, enhancing the flexibility of robotic actions. [384] proposed a vision-based full-body control framework. By connecting a robotic arm and a robotic dog, utilizing all degrees of freedom (12 joints in legs, 6 joints in the arm, and 1 in the gripper), it tracks the robot dog’s speed and the robotic arm’s end-effector position, achieving more flexible control. Some works [385, 386] employed traditional methods to control bipedal robot walking. MIT’s Cheetah 3 [387], ANYmal [388], and Atlas [389] used robust walking controllers to manage the robots. These robots can be used for more agile motion tasks, such as jumping or overcoming various obstacles [390, 391, 392, 393, 394]. Other works [395, 396] focused on the control of humanoid robots, to perform various actions like humans and mimic human behaviors. Fig .18 illustrates some examples.
@@ VIII Challenges and Future Directions
For embodied tasks, it is necessary to introduce causal representation learning. This involves establishing spatial-temporal causal relations across modalities through interactive instructions and state predictions, forming a representation learning system based on interaction and deduction. Moreover, agents need to understand the affordances of objects to achieve adaptive task planning and long-distance autonomous navigation in dynamic scenes. To optimize decision-making, it is necessary to combine counterfactual and causal intervention strategies [400], trace causality from counterfactual and causal intervention perspectives, reduce exploration iterations, and optimize decisions. Constructing a causal graph based on world knowledge and driving sim-to-real transfer of agents through active causal reasoning will form a unified framework for embodied perception, reasoning, and interaction.